### v1

@@
[…]
, runauthor = F. A. Author, S. B. Author, T. C. Author, abstract = Large-scale simulations on supercomputers have become important tools for users. However, their scalability remains a problem due to the huge communication cost among parallel processes. Most of the existing communication latency analysis methods rely on the physical link layer information, which is only available to administrators. In this paper, a framework called PCLVis is proposed to help general users analyze process communication latency (PCL) events. Instead of the physical link layer information, the PCLVis uses the MPI process communication data for the analysis. First, a spatial PCL event locating method is developed. All processes with high correlation are classified into a single cluster by constructing a process-correlation tree. Second, the propagation path of PCL events is analyzed by constructing a communication-dependency-based directed acyclic graph (DAG), which can help users interactively explore a PCL event from the temporal evolution of a located PCL events cluster. In this graph, a sliding window algorithm is designed to generate the PCL events abstraction. Meanwhile, a new glyph called communication state glyph (CS-Glyph) is designed for each process to show its communication states, including its in/out messages and load balance. Each leaf node can be further unfolded to view additional information. Third, a PCL event attribution strategy is formulated to help users optimize their simulations. The effectiveness of the PCLVis framework is demonstrated by analyzing the PCL events of several simulations running on the TH-1A supercomputer. By using the proposed framework, users can greatly improve the efficiency of their simulations. , keywords = PCLVis, communication latency, large-scale simulation, visual analytics, copyright = The Author(s),
111$\quad$1 College of Intelligence and Computing, Tianjin University, Tianjin, 300072, China. E-mail: bichongke@tju.edu.cn, gao_xin_private@163.com, fubaofeng96@163.com.
222$\quad$2 School of Data Science, Fudan University, Shanghai, 200433, China. E-mail: 22110980020@m.fudan.edu.cn, simingchen@fudan.edu.cn.
333$\quad$3 School of Computer Science and Engineering, Central South University, Changsha, 410083, China. E-mail: zhaoying@csu.edu.cn.
444$\quad$4 School of Computer Science and Technology, Shandong University, Qingdao, 266237, China. E-mail: cloudseawang@gmail.com.
Manuscript received: xxxx-xx-xx; accepted: xxxx-xx-xx
@@ 1 Introduction
However, the performance of simulation applications can be curtailed by bottlenecks in process communication[1, 2]. Communication delay events can unpredictably manifest at any juncture, prompting an imperative to dissect the underlying causes. Existing methods for communication latency analysis predominantly lean on the physical link layer, furnishing granular information encompassing start and end points, routing paths, and intermediate waypoints (routers). This information empowers the accurate identification of communication latency events. Nonetheless, the accessibility of physical link layer insights is confined to system administrators, limiting general users’ access.
@@
To bridge this gap, we introduce ’PCLVis,’ an innovative framework tailored for the wider supercomputer user base. PCLVis facilitates the visual analysis of process communication latency (PCL) events utilizing readily available MPI process communication data. First, a space-PCL event localization method is developed to localize processes involved in PCL events. All processes with a high correlation are classified into a single cluster by constructing a process-correlation tree. Second, the propagation path of PCL events is analyzed by constructing a communication-dependency-based directed acyclic graph (DAG), which can help users interactively explore a PCL event from the temporal evolution of a located PCL events cluster. In the end, A PCL events attribution strategy is also designed to help users optimize their simulations. And in summary, the graphics user interface of PCLVis is shown in Figure 1.
@@ 4.1 Criteria for Defining Communication Latency Within a Region
In the following tasks, we define two latency criteria (transmission time) for intra-node and inter-node communication. The difference between these criteria lies in whether two communication processes are in the same physical nodes or not given that the transmission speed of inter-node communication is much lower than that of intra-node communication [65].
@@
Where $t_{msg}$ is the real transmission time of $msg$, and $C_{msg}$ is the latency criterion of $msg$. If $L_{msg}>1$, then $msg$ is considered delayed. A larger $L_{msg}$ corresponds to a longer delay. A high PCL event is defined as the communication event with the larger $L_{msg}$.
@@ 4.2.1 Clustering Method
    Taking $-\log(\cdot)$ on both sides, we obtain the additive inequality:
@@ 4.2.2 Visualization of Communication Regions
To improve the visual experience of users, we use 2D force-directed edge building (FDEB), an edge bundling algorithm, to avoid cluttered connections. The system also provides many functions, including zooming in/out, translating, rotating, and dragging, for users to further observe and analyze the clustering results.
@@ 5.1 Temporal Evolution of Process Communication Latency Events Abstraction
First, $Ave_{region}$ is defined as the average latency of all messages in the $region$ (Section 4.1) . Second, we define two important features to be extracted, namely, the growth trend period and steady trend period, as illustrated in Figure 5. For other unimportant periods, we reserve three timestamps to compress the time period, namely, start, mid, and end. Based on the features defined above, we use the sliding-window-based algorithm to generate the temporal latency abstraction.
@@
Figure 6: Evolution of the latency of a communication region. The abstraction of communication latency of segments is generated based on temporal latency features. A slider component is used to select a certain period to analyze.
@@ 5.2 Communication Dependency Among Processes
Based on the physical time order of communication events, the logical time is calculated using the Vector Clock approach [68]. This approach involves the following mathematical formulation:
Based on the physical time order of communication events, the logical time is calculated using the Vector Clock approach [68]. This approach involves the following mathematical formulation:
@@
Figure 7: Calculating the logical time using Vector Clock. Three processes are involved. The x-axis represents the time, whereas the blue line represents the process of communication.
@@ 5.3.1 DAG Construction
      if $ev$ is a Send event then
@@
                find the newest created node $v_{i}$ (there may exist multiple nodes in $V$ which satisfy the condition) such that $v_{i}.pid=ev.pid$, push $ev$ into $v_{i}.eventlist$;
@@
                               push $ev$ into $v_{i}.eventlist$;
@@
                                   if the Send event corresponding to $ev$ is in $v_{j}.eventlist$ then
@@ 5.3.2 Visual Designs in DAG
    Overview and Zooming: Users can effortlessly access the communication status information for the entire DAG by employing familiar actions like translating and zooming in/out. These interactions facilitate a holistic understanding of communication patterns and statuses across the system.
@@ 7.1 MiniFE
Figure 11(a) shows the 5 communication regions of MiniFE. The processes within each of these regions communicate with one another intensively. We select the yellow region from Figure 11(a1), which has many PCL events. Figure 11(b) shows the PCL temporal evolution of the yellow region. The color of durations #1 and #2 is unusually red, which confirms the presence of communication latency in these durations. Our system only has one DAG for users at the same time. Nevertheless, to better display the effect and save space, Figure 11(c) flips the DAG for display and places the DAGs of durations #1 and duration #2 together.
Analysis of Duration #1:
To analyze the causes of PCL events in duration #1, we display the communication details of the process inside the yellow region during duration #1 in Figure 11(c-duration #1). The upper CS-Glyphs represent the sender, whereas the lower CS-Glyphs represent the receiver. The bar chart in each CS-Glyph is colored red, thereby indicating that a significant amount of time has been spent on communication. These CS-Glyphs have the same characteristics. Specifically, the ellipses of CS-Glyph pairs are flat and have different colors, which suggests that those processes belonging to different nodes communicate with one another intensively. We combine with Figure 11(d) for in-depth analysis. In Figure 11(d1), the proportion of inter-node communication is larger than that of intra-node communication, and the number of internode communications in duration #1 rapidly increases. The message transmission time is affected by background traffic when the messages need to be transmitted across inter-node links. Figure 11(d3) shows that the transmission time increases dramatically in duration#1, whereas Figure 11(d2) shows that the communication load balance of MiniFE has no significant impact on performance. Therefore, the PCL events in duration #1 are caused by poor process-to-processor mapping.
@@
Analysis of Duration #2:
Figure 11(c-duration#2) shows the communication dependency and state of PCL events in duration #2. Compared with Figure 11(c-duration#1), this view has fewer links, and the height of the bar charts is higher. Figure 11(d) shows that the number of communication decreases (Figure 11(d1)), but the transmission time of large messages greatly increases to 47462 bytes (Figure 11(d3)), hence suggesting that the transmission of large bytes in MiniFE leads to communication latency. Therefore, the PCL events in duration #2 are caused by background traffic. When the program is disturbed by background traffic, the system will prompt users to change the job node or adjust the simulation running time.
@@ 9 Conclusion
The NAS parallel benchmarks conjugate gradient and MiniFE software packages used in this study are both publicly available under open source licenses.
The NAS parallel benchmarks can be downloaded from the following website: [https://www.nas.nasa.gov/software/npb.html].
MiniFE can be downloaded from the following website: [https://computation.llnl.gov/projects/minife/].
Declaration of competing interest
The authors have no competing interests to declare that are relevant to the content of this article.
Acknowledgements
This work was partially supported by the National Key R&D Program of China under Grand No. 2021YFE0108400, and partly supported by the National Natural Science Foundation of China under Grant No. 62172294.
@@ Acknowledgements
Chongke Bi, Xin Gao, Baofeng Fu, Yuheng Zhao, Siming Chen, Ying Zhao, Yunhai Wang contributed to this paper in the following ways:
Chongke Bi: led and supported the design and implementation of the research, provided funding and equipment support for the research, and at the same time guided and revised the writing of the paper, and put forward important opinions and suggestions.
Xin Gao: participated in the design and implementation of the research, the writing and revision of the paper, and the later maintenance of the program.
Baofeng Fu: responsible for the implementation and maintenance of the research, participated in the writing and modification of the paper, and the maintenance of the program later.
Yuheng Zhao: participated in the experimental design, data collection, and made major contributions to the later debugging of the program.
Siming Chen: made important contributions to the later stage debugging of the program, and made important comments and suggestions on the guidance and revision of the manuscript writing.
Ying Zhao: participated in the research design, guided and revised the manuscript writing and made important comments and suggestions.
Yunhai Wang: guide and revise the manuscript writing, put forward important opinions and suggestions, and guide the author to complete the writing and publication of the paper.
Author biography
{biography}
[18]Chongke Bi received the B.Sc. (Eng.) degree and the M.Sc. (Eng.) degree from Shandong University, China, in 2004 and 2007, respectively, and the Ph.D. (Sci.) degree from the University of Tokyo, Japan, in 2012. After that, as a researcher in RIKEN, Japan, he was focused on research in the field of visual analysis of big data on supercomputer from 2012 to 2016. He is currently an associate Professor at Tianjin University. His current research interests include visualization and high performance computing.
{biography}
[19]Xin Gao received the B.Sc. (Sci.) degree from Qufu Normal University, China, in 2019, and the M.Sc. (Eng.) degree from Ludong University, China, in 2023. He is currently pursuing a Ph.D. (Eng.) degree from the Tianjin University, China. His current research interests include visualization and high performance computing.
{biography}
[20]Baofeng Fu received the B.Sc. (Eng.) degree from Qingdao University, China, in 2019, and the M.Sc. (Eng.) degree from Tianjin University, China, in 2022. Here research interests include visualization and high performance computing.
{biography}
[21]Yuheng Zhao Yuheng Zhao is currently a Ph.D student at the School of Data Science, Fudan University. Her main research interests include social media visualization and text visual analytics.
{biography}
[22]Siming Chen received the Ph.D. degree in computer and information technology from Peking University, Beijing, China, in 2017. He is currently an Associate Professor with the School of Data Science, Fudan University, Shanghai, China. Prior to this, he was a Research Scientist with Fraunhofer Institute IAIS and a Postdoc Researcher with the University of Bonn, Bonn, Germany. He has authored or coauthored more than 30 papers in visualization field, including more than 10 IEEE VIS/TVCG/EuroVis papers. His research interests include general visual analytics.
{biography}
[23]Ying Zhao is a Professor at the School of Computer Science and Engineering, Central South University, China. He received his Ph. D. degree in Computer Application Technology from Central South University in 2014. His research interests include visualization and visual analytics.
{biography}
[24]Yunhai Wang is a professor in the School of Computer Science and Technology at the Shandong University Qingdao campus. He leads the Interactive Data Exploration System (IDEAS) lab that aims to enhance people’s ability to understand and communicate data through the design of automated visualization and visual analytics systems. His research focuses on the question of how can we automatically design an effective visualization that is best suited to pursue a given task on given input data.

### v2

@@
Abstract
Large-scale simulations on supercomputers have become important tools for users. However, their scalability remains a problem due to the huge communication cost among parallel processes. Most of the existing communication latency analysis methods rely on the physical link layer information, which is only available to administrators. In this paper, a framework called PCLVis is proposed to help general users analyze process communication latency (PCL) events. Instead of the physical link layer information, the PCLVis uses the MPI process communication data for the analysis. First, a spatial PCL event locating method is developed. All processes with high correlation are classified into a single cluster by constructing a process-correlation tree. Second, the propagation path of PCL events is analyzed by constructing a communication-dependency-based directed acyclic graph (DAG), which can help users interactively explore a PCL event from the temporal evolution of a located PCL events cluster. In this graph, a sliding window algorithm is designed to generate the PCL events abstraction. Meanwhile, a new glyph called communication state glyph (CS-Glyph) is designed for each process to show its communication states, including its I/o messages and load balance. Each leaf node can be further unfolded to view additional information. Third, a PCL event attribution strategy is formulated to help users optimize their simulations. The effectiveness of the PCLVis framework is demonstrated by analyzing the PCL events of several simulations running on the TH-1A supercomputer. By using the proposed framework, users can greatly improve the efficiency of their simulations.
keywords:
PCLVis, communication latency, large-scale simulation, visual analytics
@@ 1 Introduction
However, the performance of simulation applications can be curtailed by bottlenecks in process communication[1]. Communication delay events can unpredictably manifest at any juncture, prompting an imperative to dissect the underlying causes. Existing methods for communication latency analysis predominantly lean on the physical link layer, furnishing granular information encompassing start and end points, routing paths, and intermediate waypoints (routers). This information empowers the accurate identification of communication latency events. Nonetheless, the accessibility of physical link layer insights is confined to system administrators, limiting general users’ access.
@@
To bridge this gap, we introduce ’PCLVis’, an innovative framework tailored for the wider supercomputer user base. PCLVis facilitates the visual analysis of process communication latency (PCL) events utilizing readily available MPI process communication data. First, a space-PCL event localization method is developed to localize processes involved in PCL events. All processes with a high correlation are classified into a single cluster by constructing a process-correlation tree. Second, the propagation path of PCL events is analyzed by constructing a communication-dependency-based directed acyclic graph (DAG), which can help users interactively explore a PCL event from the temporal evolution of a located PCL events cluster. In the end, A PCL events attribution strategy is also designed to help users optimize their simulations. And in summary, the graphics user interface of PCLVis is shown in Figure 1.
@@ 4.1 Criteria for Defining Communication Latency Within a Region
In the following tasks, we define two latency criteria (transmission time) for intra-node and inter-node communication. The difference between these criteria lies in whether two communication processes are in the same physical nodes or not given that the transmission speed of inter-node communication is much lower than that of intra-node communication.
@@
Where $t_{msg}$ is the real transmission time of msgmsg, and $C_{msg}$ is the latency criterion of msgmsg. If $L_{msg}>1$, then msgmsg is considered delayed. A larger $L_{msg}$ corresponds to a longer delay. A high PCL event is defined as the communication event with the larger $L_{msg}$.
@@ 4.2.1 Clustering Method
    Taking $\log(\cdot)$ on both sides, we obtain the additive inequality:
@@ 4.2.2 Visualization of Communication Regions
To improve the visual experience of users, we use 2D force-directed edge building (FDEB), an edge bundling algorithm, to avoid cluttered connections. The system also provides many functions, including zooming I/O, translating, rotating, and dragging, for users to further observe and analyze the clustering results.
@@ 5.1 Temporal Evolution of Process Communication Latency Events Abstraction
First, $Ave_{region}$ is defined as the average latency of all messages in the regionregion (Section 4.1) . Second, we define two important features to be extracted, namely, the growth trend period and steady trend period, as illustrated in Figure 5. For other unimportant periods, we reserve three timestamps to compress the time period, namely, start, mid, and end. Based on the features defined above, we use the sliding-window-based algorithm to generate the temporal latency abstraction.
@@
Figure 6: Evolution of the latency of a communication region. The abstraction of the communication latency of segments is generated based on temporal latency features. A slider component is used to select a certain period to analyze.
@@ 5.2 Communication Dependency Among Processes
Based on the physical time order of communication events, the logical time is calculated using the Vector Clock approach. This approach involves the following mathematical formulation:
Based on the physical time order of communication events, the logical time is calculated using the Vector Clock approach. This approach involves the following mathematical formulation:
@@
Figure 7: Calculating the logical time using Vector Clock. Three processes are involved. The x-axis represents time, whereas the blue line represents the process of communication.
@@ 5.3.1 DAG Construction
 if evev is a Send event then
@@
         find the newest created node $v_{i}$ (there may exist multiple nodes in $V$ which satisfy the condition) such that $v_{i}.pid=ev.pid$, push evev into $v_{i}.eventlist$;
@@
         push evev into $v_{i}.eventlist$;
@@
    if the Send event corresponding to evev is in $v_{j}.eventlist$ then
@@ 5.3.2 Visual Designs in DAG
    Overview and Zooming: Users can effortlessly access the communication status information for the entire DAG by employing familiar actions like translating and zooming I/O. These interactions facilitate a holistic understanding of communication patterns and statuses across the system.
@@ 7.1 MiniFE
Figure 11(a) shows the 5 communication regions of MiniFE. The processes within each of these regions communicate with one another intensively. We select the yellow region from Figure 11(a1), which has many PCL events. Figure 11(b) shows the PCL temporal evolution of the yellow region. The color of durations 1 and 2 is unusually red, which confirms the presence of communication latency in these durations. Our system only has one DAG for users at the same time. Nevertheless, to better display the effect and save space, Figure 11(c) flips the DAG for display and places the DAGs of duration 1 and duration 2 together.
Analysis of Duration 1:
To analyze the causes of PCL events in duration 1, we display the communication details of the process inside the yellow region during duration 1 in Figure 11(c-duration 1). The upper CS-Glyphs represent the sender, whereas the lower CS-Glyphs represent the receiver. The bar chart in each CS-Glyph is colored red, thereby indicating that a significant amount of time has been spent on communication. These CS-Glyphs have the same characteristics. Specifically, the ellipses of CS-Glyph pairs are flat and have different colors, which suggests that those processes belonging to different nodes communicate with one another intensively. We combine with Figure 11(d) for in-depth analysis. In Figure 11(d1), the proportion of inter-node communication is larger than that of intra-node communication, and the number of internode communications in duration 1 rapidly increases. The message transmission time is affected by background traffic when the messages need to be transmitted across inter-node links. Figure 11(d3) shows that the transmission time increases dramatically in duration 1, whereas Figure 11(d2) shows that the communication load balance of MiniFE has no significant impact on performance. Therefore, the PCL events in duration 1 are caused by poor process-to-processor mapping.
@@
Analysis of Duration 2:
Figure 11(c-duration 2) shows the communication dependency and state of PCL events in duration 2. Compared with Figure 11(c-duration 1), this view has fewer links, and the height of the bar charts is higher. Figure 11(d) shows that the number of communication decreases (Figure 11(d1)), but the transmission time of large messages greatly increases to 47462 bytes (Figure 11(d3)), hence suggesting that the transmission of large bytes in MiniFE leads to communication latency. Therefore, the PCL events in duration 2 are caused by background traffic. When the program is disturbed by background traffic, the system will prompt users to change the job node or adjust the simulation running time.
@@ 9 Conclusion
Acknowledgments
This work was funded mainly by the National Natural Science Foundation of China under Grant No. 62172294, the Tianjin Natural Science Foundation under Grant No. 23JCYBJC00440, and the Tianjin Science and Technology Project under Grant No. 24YDTPJC00400.
@@ Acknowledgments
The MPI communication trace datasets used in this study were generated from large-scale simulation runs on the supercomputer. During the simulations, we collected raw parallel execution traces using the Tuning Analysis Utility (TAU). These raw traces were subsequently preprocessed to extract relevant communication events, including MPI rank, event type, timestamp, source/destination ranks, and message size. The processed datasets were then used for the spatial latency localization, temporal abstraction, and communication-dependency analysis presented in this paper.
Competing interests
Not applicable.
@@
Chongke Bi, Xin Gao, and Baofeng Fu designed and implemented the PCLVis framework and developed the algorithms for spatial latency localization, temporal abstraction, and communication-dependency DAG construction. Yuheng Zhao and Siming Chen analyzed and processed the MPI communication trace data from the simulations. Ying Zhao and Lu Yang contributed to the latency attribution strategy, experimental design, and overall supervision of the research.
Funding
This work was funded mainly by the National Natural Science Foundation of China under Grant No. 62172294, the Tianjin Natural Science Foundation under Grant No. 23JCYBJC00440, and the Tianjin Science and Technology Project under Grant No. 24YDTPJC00400.